OpenStudy (1645323):

The function y = -2.38x2 + 45.24x + 800 models the population of a local high school and x = 0 represents the year 1985. Use the Quadratic Formula to estimate the years in which 1000 students attended this high school? Select one: a. 2017 and 2032 b. 2010 and 2019 c. 1990 and 2010 d. 1992 and 2017

OpenStudy (1645323):

@maths911

OpenStudy (1645323):

1992 is right but not 2017

OpenStudy (jdoe0001):

\(\large \begin{array}{rrll} y = &-2.38x^2 + 45.24x + 800 \\ \quad \\ {\color{red}{ \textit{population of a local high school}}}=&-2.38x^2 + 45.24x + 800 \end{array} \\ \quad \\ \bf {\color{red}{ 1000}}=-2.38x^2 + 45.24x + 800\implies {\color{blue}{ 2.38}}x^2{\color{red}{ -45.24}}x{\color{green}{ +200}} \\ \quad \\ x= \cfrac{ - {\color{red}{ b}} \pm \sqrt { {\color{red}{ b}}^2 -4{\color{blue}{ a}}{\color{green}{ c}}}}{2{\color{blue}{ a}}}\)

OpenStudy (jdoe0001):

hmm kinda forgot the 0... well \( \begin{array}{rrll} y = &-2.38x^2 + 45.24x + 800 \\ \quad \\ {\color{red}{ \textit{population of a local high school}}}=&-2.38x^2 + 45.24x + 800 \end{array} \\ \quad \\\\ \quad \\ \bf {\color{red}{ 1000}}=-2.38x^2 + 45.24x + 800\implies {\color{blue}{ 2.38}}x^2{\color{red}{ -45.24}}x{\color{green}{ +200}}=0 \\ \quad \\ x= \cfrac{ - {\color{red}{ b}} \pm \sqrt { {\color{red}{ b}}^2 -4{\color{blue}{ a}}{\color{green}{ c}}}}{2{\color{blue}{ a}}}\)

OpenStudy (1645323):

so is it c?

OpenStudy (jdoe0001):

dunno... what did you get for the discriminant?

OpenStudy (1645323):

9662.6576

OpenStudy (1645323):

@jdoe0001

OpenStudy (jdoe0001):

well... you can just plug in the values to see what "t" is